the 600N force applied to the bracket at A is to be replaced by two forces , Fa in the a-a direction and Fb in the b-b direction, which together produce the same effect on the bracket as that of 600- N force determine Fa and Fb

1
Expert's answer
2022-01-10T09:15:40-0500

"F_a=\frac F{\cos 30°}=\frac{2\cdot 600}{\sqrt3}=692.82~N."

"F_b=\frac F{2\sin60°}=\frac{600}{\sqrt3}=346.41~N."